You own the following stocks in your portfolio. What is the beta of your portfolio.

Stock Invested Amount Beta
A 9,289 1.36
B 4,944 0.59
C 7,824 1.12
D 3,079 0.81
Answer #1

Total amount invested=9,289+4,944+7,824+3,079=$25136

Portfolio beta=Respective beta*Respective weight

=(9,289/25136*1.36)+(4,944/25136*0.59)+(7,824/25136*1.12)+(3,079/25136*0.81)

=1.07(Approx).